OpenStudy (anonymous):

A certain species of tree grows an average of 3.1 cm per week. Write an equation for the sequence that represents the weekly height of this tree in centimeters if the measurements begin when the tree is 600 centimeters tall. Is it h = 600 + 3.1w or can you simplify it?

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

no need for simplification

OpenStudy (anonymous):

h = 600 + 3.1w n = a + (n-1) d n = 600 + (n-1) 3.1 n = 600 + 3.1n -3.1 n = 596.9 + 3.1n

OpenStudy (anonymous):

So it's right either way?